OpenStudy (anonymous):

The cost of craft beads varies directly with the number purchased. The cost of 25 beads is $10. How much does 8 beads cost? Write an equation relating cost and beads purchased Use the equation to find the number of 8 beads purchased.

OpenStudy (anonymous):

\[\frac{10}{25}=\frac{x}{8}\iff x=\frac{8\times 10}{25}\]